OpenStudy (anonymous):

Solve: 4(2x − 1) − 3 = 9(x − 4)

OpenStudy (anonymous):

4(2x - 1) - 3 = 9(x - 4) 8x - 4 - 3 = 9x - 36... expand out the brackets 8x - 7 = 9x - 36... collect like terms -x - 7 = -36... move the 9x over to the left -x = -29... move the -7 to the right x = 29... divide by -1 x = 29

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

welcome :P